OpenStudy (anonymous):

Simplify. x^2−5x+4/x−1

OpenStudy (anonymous):

\[x^2-5x+4/x-1\]

OpenStudy (anonymous):

we did LOL!

OpenStudy (anonymous):

it wasnt clear

OpenStudy (anonymous):

oh okay then

OpenStudy (anonymous):

you can factor the numerator and you'll get:\[=(x-1)(x-4)/(x-4)\] then you can cancel what you have common from the top and bottom :) and you'll get: \[= (x-1)\] is it clear now? ^_^

OpenStudy (anonymous):

wait, you'll get (x-4) lol sorry!

OpenStudy (anonymous):

it's : (x-1)(x-4)/(x-1) = (x-4) sorry for the mess :)

OpenStudy (anonymous):

are you sure its not (x^2) -(5x) +(4/x) - (1) lol hahahah jsut making sure xD

OpenStudy (anonymous):

is it clear now JAPAN? :)

OpenStudy (anonymous):

i'm sure

OpenStudy (anonymous):

and yes looks good to me

OpenStudy (anonymous):

that's good ^_^

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

X^2- 5x + 4/x-1 U can write X^2- 5x + 4 as (x-1) . (x-4), then (x-1) . (x-4)/(x-1) = x-4.... sstarica is true japan,,,, its simple question,...

OpenStudy (anonymous):

it's alright, even the easiest questions may look tough :)

OpenStudy (anonymous):

well i'm a idiot who spends his time in literature

OpenStudy (anonymous):

no, don't say that , it's not true.Look at me for instance, I do silly mistakes such as instead adding I subtract and the vice versa, but then I practiced to stop such things. You'r not an idiot, you just need mroe practice ^_^ and I wish you the best of luck!

OpenStudy (anonymous):

instead of*